OpenStudy (anonymous):

Write an equation of a line that passes through the point (2, 3) and is parallel to the line y =3/2 x + 5. y = x y = x - 7 y = x y = x + 7

OpenStudy (anonymous):

y = -2/3 x y =-2/3 x - 7 y =3/2 x y =3/2 x + 7

OpenStudy (anonymous):

first you have to get the slope of the line ? any ideas?

OpenStudy (anonymous):

no :/

OpenStudy (anonymous):

if you have a two parallel lines their slopes are the same the slope of the line =?

OpenStudy (anonymous):

the slope is the same so its 3/2

OpenStudy (anonymous):

right now you have a slope and a point and the equation of the line in this case \[(y -y _{1})=m(x _-x _{1})\]

OpenStudy (anonymous):

where m=slope (x1,y1)=point

OpenStudy (anonymous):

(y-3)=3/2(x-2)

OpenStudy (anonymous):

eq. of any line parallel to y=3/2 x +5 is y=3/2 x+c if it passes through (2,3) find c and then the eq.

OpenStudy (anonymous):

right then simplify the equation to y=(3/2)x-3+3 y=(3/2)x

OpenStudy (anonymous):

once you work it out you get 1 right ? so the answer is c ?

OpenStudy (anonymous):

yay i get it now!! thank you guys!!

OpenStudy (anonymous):

anytime :)